1- Find electric field intensity due to dipole of dipole moment 4.4x10-19C-m at a point 30nm away along bisector axis?

2- A point charge of 1x 10-6C is at centre of Gaussian surface with cubical shape 1m on edge. Find flux through surface? 

3- Consider a point charge 2C, what is the radius of equipotential surface having potential 20V? 

1
Expert's answer
2020-12-31T08:06:04-0500

Answer

1-electric field due to dipole is

"E=\frac{kp}{r^3}"

= "\frac{9\times10^9\times4.4\times10^{-19}}{27\times 10^{-24} }"

"=1.46\times10^{14}V/m"

2-flux through all surface

"\phi=\frac{q}{\epsilon_0}=\frac{10^{-6}}{8.85\times{10^{-12}}}=11.30\times{10^{4}}"V-m


3-radius of equipotential surface is given by

"r=\frac{kq}{V}=\frac{9\times10^9\times2}{20}=9\times10^8m"